OpenStudy (anonymous):

solve using the addition methd 2x-3y=8 and 3x+2y=-1

OpenStudy (anonymous):

help.......

OpenStudy (anonymous):

Multiply the first equation by \(3\) and the second one by \(2\) and then add them up, the y term will be cancelled. It will be easy then to solve for x. Finally use that value of x into one of the equations and get y.

OpenStudy (anonymous):

The first by \(2\) and the second by \(3\) :D

OpenStudy (amistre64):

2x-3y=8 [-3/2] 3x+2y=-1 -3x +9/2 y = -12 3x +2 y = -1 ---------------- 13/2 y = -13 y = -2

OpenStudy (amistre64):

3x+2y=-1 3x+2(-2)=-1 3x -4 = -1 3x = 3; x=1

OpenStudy (anonymous):

ty i gave u guys a medal

OpenStudy (amistre64):

:) youre welcome